OpenStudy (anonymous):

Can someone please give me the derivation of Newtons Law of Universal Gravitation

OpenStudy (ujjwal):

\[F \alpha 1\div r ^{2}\]\[F \alpha m1*m2\] from above two relations \[F \alpha \left( m1*m2 \right)\div r ^{2}\] or, \[F= G \left( m1*m2 \right)\div r ^{2}\] where G is universal gravitational constant m1= mass of 1st body m2= mass of 2nd body r= distance of separarion betweeen them

OpenStudy (anonymous):

Every body in universe attracts every other body with a force which is directly proportional to the product of their masses and inversely proportional to the square of the distance b/w their centers. \[F \alpha m1.m2\] \[F \alpha1/r2\] Combining these two Relation:

OpenStudy (anonymous):

u cant derive newtons law of gravitations ! can u ?? its a LAW !! :/ BUT U can tell its directly propotional to mass and inversely to square of the distance ... thats all

OpenStudy (anonymous):

It is the apple and the moon problem. http://learner.org/vod/vod_window.html?pid=555